in a rectangle the ratio of the length to width is 5:2, and its perimeter is 126 centimeter. Find the width of the rectangle?
A. 9cm B. 18 cm C.45cm D.50.4cm

Answer by checkley79(3341)

5:2 HAS 7 ELEMENTS.
126/7=18
MULTIPLY EACH ELEMENT BY 18
5*18:2*18=90:36
THUS THE SIDES ARE:
2L=90
L=90/2
L=45
2W=36
W=36/2
W=18
PROOF:
126=2L+2w
126=90+36
126=126